# WorldLLM: Improving LLMs’ world modeling using curiosity-driven theory-making

### v1

@@ 1 Introduction
Large Language Models (LLMs) possess broad knowledge about the world, but leveraging this knowledge for precise dynamics modeling remains challenging (Vafa et al., 2024). While LLMs can engage in general reasoning, they struggle to make accurate predictions in specific domains with structured observations and dynamics, such as physics simulations or video games. This limitation stems from the gap between their general capabilities and the need for grounded, domain-specific understanding (Carta et al., 2023; Cross et al., 2025; Bender & Koller, 2020).
@@
The challenge of learning world models has been extensively studied in model-based reinforcement learning (RL), as depicted in Moerland et al. (2023). Traditional approaches focus on learning forward dynamics models through direct interactions, typically using neural networks to predict next observations given current states and actions. Recent work has shown promise in using LLMs for world modeling by fine-tuning them on embodied experiences (Xiang et al., 2023). However, such an approach is particularly computationally costly given the large amount of interaction data required as well as the cost of gradient-based finetuning of LLMs. Contrary to this, LLMs are known to possess general knowledge that can be adapted using careful prompting (Zhang et al., 2024).
@@
In a parallel line of work, theory-based RL takes inspiration from humans’ inductive reasoning and intuitive theory-making to suggest focusing on inferring structured and compositional world models (e.g. programs) (Tsividis et al., 2021). Current approaches either generate programs to capture environment dynamics directly (Tsividis et al., 2021; Tang et al., 2024a; Li et al., 2024) or use natural language rules as an intermediate representation (Piriyakulkij et al., 2024; Wang et al., 2024). Leveraging the world model’s structure as well as human-provided prior distributions over possible world models, theory-based methods are usually more sample-efficient than gradient-based ones. More importantly, the resulting world model often shows stronger generalization abilities. However, contrary to gradient-based approaches that usually do not require any expert-given knowledge, theory-based RL methods mostly rely on hand-crafted theory spaces. Finally, one major challenge in learning a world model lies in the data collection. Indeed, collecting transitions that cover the space of possible outcomes in the environment is key for learning an accurate world model. Notably, both classic model-based RL approach and theory-based RL methods agree on the importance of active data collection focusing on evidence with low likelihood under the current model (Schmidhuber, 1991a; Pathak et al., 2017; Burda et al., 2019b; Piriyakulkij et al., 2024).
@@
In this paper, we present WorldLLM, a framework for autonomous improvement of an LLM’s world modeling abilities. Our approach combines 1) probabilistic theory induction to produce hypotheses that are given in our LLM’s prompt to improve its predictions and 2) curiosity-driven RL to explore the environment and collect transitions poorly predicted with the current hypotheses. Formally, our LLM’s world model is the conditional probability $P(s_{t+1}|s_{t},a_{t},H)$, where $s_{t}$ represents a state, $a_{t}$ an action, and $H$ a set of natural language hypothesized theories. This probability is computed by the LLM by giving it $s_{t}$, $a_{t}$ and $H$ in its prompt and taking the probability of $s_{t+1}$ to follow this prompt. Our key insight is that natural language theories can help ground an LLM’s broad knowledge into precise predictive power by providing domain-specific rules (Zhang et al., 2024). Our approach consists of three interacting components: (1) our LLM that computes $P(s_{t+1}|s_{t},a_{t},H)$ by conditioning its predictions on both a state-action pair and the current hypotheses, (2) a theory generator that updates natural language hypotheses using Bayesian inference, and (3) a curiosity-driven reinforcement learning agent trained to collect evidence against the current hypotheses. Inspired by how humans, from children to scientists (Piaget, 1954; Gopnik, 1996), actively update their internal world model by performing experiments, our agent’s exploration provides new evidence for hypothesis refinement, creating a virtuous cycle of improvement.
@@
We demonstrate our approach in a video game environment where agents manipulate and combine objects, showing that WorldLLM successfully learns accurate predictive models while generating human-interpretable theories about environment dynamics. This work contributes to a growing body of research on improving LLMs’ world modeling capabilities and grounding their knowledge in specific domains (Carta et al., 2023; Xiang et al., 2023; Patel et al., 2020). By combining ideas from theory-based RL, Bayesian inference, and active exploration, we provide a framework for learning structured, interpretable world models that leverage both the broad knowledge of LLMs and domain-specific experiences without any costly gradient-based learning.
@@ 2.1 Problem statement
We consider an environment framed as a Markov Decision Process $\mathcal{M=(S,A,R,T)}$ with $\mathcal{S}$ the state space, $\mathcal{A}$ the action space, $\mathcal{R:S\times A\times S}\mapsto\mathbb{R}$ the reward function and $\mathcal{T:S\times A\mapsto S}$ the transition function. In particular, as our main objective is to study world modeling abilities, we consider reward-less environments where no particular task must be solved, i.e. the reward function always returns 0. We focus on textual environments where, given a vocabulary $\mathcal{V}$ and a maximum sentence length $N$, observations $o\in\mathcal{S}\subset\mathcal{V}^{N}$ and actions $a\in\mathcal{A}\subset\mathcal{V}^{N}$ are represented in natural language.
In such a setting, we study how to rapidly learn an LLM-based parametric model $\mathbb{P}^{LLM}$ of the transition function $\mathcal{T}$ (also called forward model) given a budget of experiences $T$ allowed to an agent. This parametric model outputs a probability distribution over all possible next states given a state-action pair. The natural approach from the RL literature to improve this forward model uses gradient-based updates (Ha & Schmidhuber, 2018; Hafner et al., 2020; Schrittwieser et al., 2020). However, the scale of modern LLMs makes such an approach particularly costly. LLMs’ natural in-context learning abilities open up an alternative approach: providing natural language hypotheses in the prompt to improve its predictions. This method is less costly, known as more sample efficient (Le Scao & Rush, 2021), and provides human-interpretable updates. Consequently, the parameter we seek to optimize is a set of hypotheses $\mathcal{H}\subseteq\mathcal{V}^{N}$ that, when given as input to the forward model (in addition to state-action pairs), maximizes the likelihood of the experienced next states.
We define the global optimization objective as finding the smallest set of hypothesis $H\subseteq\mathcal{H}$ that maximizes the forward model’s probability of all possible transitions in the environment:
@@ 2.2 WorldLLM
In this section, we present WorldLLM, our methodology for addressing the challenge outlined in the preceding section. Our framework comprises three interconnected components: the Statistician, the Scientist, and the Experimenter. The interactions among these components are illustrated in Figure 1. The Statistician represents our LLM-based forward model (i.e. $\mathbb{P}^{LLM}$) and is used to evaluate the current hypotheses. The set of hypotheses $H$ the Statistician uses is produced by the Scientist, another LLM. These hypotheses are derived from trajectories collected from interactions with the environment using the Experimenter. WorldLLM alternates from during $T$ iterations between evidence collection from the Experimenter and hypotheses refinement from the Scientist. In the subsequent sections, we elaborate on how these modules contribute to our objective’s optimization.
2.2.1 Statistician
We first detail the Statistician module whose role is to evaluate a set of hypotheses $H_{t}$ on collected transitions. We start by renaming, for the sake of clarity, $\mathbb{P}^{LLM}$ by $\mathbb{P}^{St}$. As described above, the Statistician is a forward model computing the likelihood $\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ of the next state $s^{\prime}$ given a state-action pair $(s,a)$ and a set of hypotheses $H_{t}$. This likelihood is obtained by the LLM by giving it $s$, $a$ and $H_{t}$ in the prompt and taking the probability of $s^{\prime}$ to follow this prompt.
2.2.2 Scientist
To improve the Statistician’s prediction as a forward model, the Scientist’s role is to search through the space of possible sets of hypotheses. To efficiently explore this space, we take inspiration from (Wang et al., 2024; Piriyakulkij et al., 2024) and use a Bayesian Inference framework. In particular, we rely on another LLM $\mathbb{P}^{Sc}$ as the proposal distribution which generates a candidate set of hypotheses following the distribution $\mathbb{P}^{Sc}(\hat{H}_{t}^{i}|\mathcal{D}_{t},\hat{H}_{t})$ given a data set of collected transitions $\mathcal{D}_{t}$ and the current best hypotheses candidate $\hat{H_{t}}$. We assume a uniform prior (i.e. no particular prior) over the space of hypotheses. For this paper, we chose the Metropolis algorithm, a variant of the Metropolis-Hastings algorithm, which iteratively generates candidate hypotheses for $n_{s}teps$. The full algorithm is displayed in Algorithm 1 and more details are provided in Appendix B.
@@ 2.2.2 Scientist
13:              Update score $current\_score=score$
@@
2.2.3 Experimenter
Given a set of hypotheses produced by the Scientist, the Experimenter must collect transitions that will help assessing and refining the hypotheses (i.e. transitions that have a low likelihood under the Statistician). We explore two approaches for our Experiment: oracle agents, which are hard-coded policies that have access to the environment’s inner dynamics, and curiosity-driven RL agents trained to collect transitions based on an intrinsic signal.
For the oracles, we propose the following policies:
@@ 2.2.3 Experimenter
Regarding our RL policies, we propose three agents trained using curiosity-based intrinsic rewards derived from $\mathbb{P}^{St}$ (Oudeyer & Kaplan, 2007). First, we introduce RL-LogP, which uses the prediction error or surprisal (Burda et al., 2019a; Pathak et al., 2017) as reward by receiving $r=-log\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ for each collected transition $(s,a,s^{\prime})$. However, this method is known to suffer from the "noisy TV problem" (Burda et al., 2019b) where unpredictable transitions give high rewards.
To address this, we inspired from methods relying on Learning Progress (Oudeyer et al., 2007; Lopes et al., 2012; Schmidhuber, 1991a) or information gain over the prediction error (Oudeyer et al., 2007; Lopes et al., 2012; Schmidhuber, 1991a) that reward transitions on which the forward model improves. We thus propose RL-ALP, which uses the absolute value of Learning Progress, allowing it to track both progress and forgetting in the forward model (Oudeyer et al., 2007; Baranes & Oudeyer, 2009):
@@
Finally, we also propose RL-ALPEXP, a variant of RL-ALP which partitions the transition space into subspaces and computes ALP over each subspace. However, the partitions are not automatically inferred (e.g. as in Baranes & Oudeyer (2013)) but hand-designed. It corresponds to a middle ground between the oracles and RL-LogP or RL-ALP as it has access to expert environment knowledge while still relying on RL. Further details are provided in Appendix A.
@@ 3.1 Experimental setup
Figure 2: Our experimental setup. We use the Playground-Text environment which features objects the agent can interact with along with textual observations and actions (a). We focus on 4 object types from Playground-Text: Water, Plants, Small Herbivores, and Big Herbivores. We show in (b) the technology tree for combining objects.
To evaluate WorldLLM’s efficiency, we use the Playground-Text environment from (Gaven et al., 2024), a textual environment which produces textual observations and receives textual commands. This environment generates a room that contains multiple objects the agent can interact with by grasping or releasing them. The agent can combine two objects, if an interaction is possible, by releasing one onto the other, as illustrated on Figure 2 b. For instance, releasing water on a carrot seed transforms them into grown carrot. In our experiments, we only consider 4 types of objects: Water, Plants (seed or grown) that require water to grow, Small Herbivores (baby or grown) that require a grown plant to grow, and Big Herbivores (baby or grown) that require 2 grown plants to grow. We evaluate WorldLLM’s ability to discover, predict and explain how to grow Plants, Small Herbivores and Big Herbivores. We provide more details on the environment in Appendix A. For analysis, we gather the possible transitions into 6 different types: Standing when moving on an object, Holding 1 when grasping an object, Holding 2 when grasping two objects at the same time, Grow Plant when releasing water on a seed, Grow S. Herbivore when releasing a grown plant on a baby small herbivore and Grow B. Herbivore when releasing two grown plants on a baby big herbivore. These types are the ones used by RL-ALPEXP.
@@
We use Phi-3-mini-4k-Instruct [^1] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct for both the Statistician and the Scientist. For our Experimenter, we either use one of our hard-coded oracles or an RL agent. Regarding oracles, O-Ideal collects transitions that match the distribution of the above test set. O-Hardest performs the optimal policy to grow the Small Herbivore then the Big Herbivore accessible. O-Curriculum plays the optimal policy to grow a Plant up to 133 iterations, then plays the optimal policy to grow the accessible Small Herbivore up to 266 iterations, and finally grows the Small Herbivore then the Big Herbivore until the end. See Appendix B.3.1 for more details on the oracles. Concerning the RL agent, its policy is a Multi-Layer Perceptron with $2$ hidden layers of $64$ units. For simplicity, this policy uses a symbolic representation of Playground-Text’s state as observation. In all experiments below, unless specified otherwise, we use $8$ different random seeds. For each seed, we perform $400$ iterations of the framework (i.e. $T$) where $150$ transitions are collected per iteration and $n_{steps}=5$ steps of the Metropolis algorithm are performed by the Scientist at each iteration. When the Experimenter is an RL agent, we collect $3600$ transitions per iteration to train it and use PPO to update the policy, while only keeping the last $150$ transitions for the Scientist to match the oracles’ hyperparameters. Our test set $\mathcal{D}_{test}$ is composed of 120 Standing, 20 Holding 1, 7 Holding 2, 12 Grow Plant, 6 Grow S. Herbivore and 3 Grow B. Herbivore.
Figure 3: Evolution of log-likelihodds computed by the Statistician (using the Scientist’s last accepted hypotheses) on the test set throughout WorldLLM’s iterations when using different oracles as Experimenters. We average per transition type over $8$ random seeds with the standard deviation represented by the error bars. We show at iteration 0 the log-likelihood from the Statistician without any hypotheses given.

### v2

@@ 1 Introduction
Large Language Models (LLMs) possess broad knowledge about the world, but leveraging this knowledge for precise dynamics modeling remains challenging (vafa_evaluating_2024). While LLMs can engage in general reasoning, they struggle to make accurate predictions in specific domains with structured observations and dynamics, such as physics simulations or video games. This limitation stems from the gap between their general capabilities and the need for grounded, domain-specific understanding (carta_grounding_2023; cross_hypothetical_2024; bender_climbing_2020).
@@
The challenge of learning world models has been extensively studied in model-based reinforcement learning (RL), as depicted in moerland_model-based_2022. Traditional approaches focus on learning forward dynamics models through direct interactions, typically using neural networks to predict next observations given current states and actions. Recent work has shown promise in using LLMs for world modeling by fine-tuning them on embodied experiences (xiang_language_2023). However, such an approach is particularly computationally costly given the large amount of interaction data required as well as the cost of gradient-based finetuning of LLMs. Contrary to this, LLMs are known to possess general knowledge that can be adapted using careful prompting (zhang_language-guided_2024).
@@
In a parallel line of work, theory-based RL takes inspiration from humans’ inductive reasoning and intuitive theory-making to suggest focusing on inferring structured and compositional world models (e.g. programs) (tsividis_human-level_2021). Current approaches either generate programs to capture environment dynamics directly (tsividis_human-level_2021; tang_worldcoder_2024; li_automated_2024) or use natural language rules as an intermediate representation (piriyakulkij_doing_2024; wang_hypothesis_2023). Leveraging the world model’s structure as well as human-provided prior distributions over possible world models, theory-based methods are usually more sample-efficient than gradient-based ones. More importantly, the resulting world model often shows stronger generalization abilities. However, contrary to gradient-based approaches that usually do not require any expert-given knowledge, theory-based RL methods mostly rely on hand-crafted theory spaces. Finally, one major challenge in learning a world model lies in the data collection. Indeed, collecting transitions that cover the space of possible outcomes in the environment is key for learning an accurate world model. Notably, both classic model-based RL approach and theory-based RL methods agree on the importance of active data collection focusing on evidence with low likelihood under the current model (schmidhuber_curious_1991; pathak_curiosity-driven_2017; burda_exploration_2018; piriyakulkij_doing_2024).
@@
In this paper, we present WorldLLM, a framework for autonomous improvement of an LLM’s world modeling abilities. Our approach combines 1) probabilistic theory induction to produce hypotheses that are given in our LLM’s prompt to improve its predictions and 2) curiosity-driven RL to explore the environment and collect transitions poorly predicted by the current hypotheses. Formally, our LLM’s world model is the conditional probability $P(s_{t+1}|s_{t},a_{t},H)$, where $s_{t}$ represents a state, $a_{t}$ an action, and $H$ a set of natural language hypothesized theories. This probability is computed by the LLM by giving it $s_{t}$, $a_{t}$, and $H$ in its prompt and taking the probability of $s_{t+1}$ to follow this prompt. Our key insight is that natural language theories can help ground an LLM’s broad knowledge into precise predictive power by providing domain-specific rules (zhang_language-guided_2024). Our approach consists of three interacting components: (1) our LLM that computes $P(s_{t+1}|s_{t},a_{t},H)$ by conditioning its predictions on both a state-action pair and the current hypotheses, (2) a theory generator that updates natural language hypotheses using Bayesian inference, and (3) a curiosity-driven reinforcement learning agent trained to collect evidence against the current hypotheses. Inspired by how humans, from children to scientists (piaget_construction_1954; gopnik_words_1997), actively update their internal world model by performing experiments, our agent’s exploration provides new evidence for hypothesis refinement, creating a virtuous cycle of improvement.
@@
We demonstrate our approach in a video game environment where agents manipulate and combine objects, showing that WorldLLM successfully learns accurate predictive models while generating human-interpretable theories about environment dynamics. This work contributes to a growing body of research on improving LLMs’ world modeling capabilities and grounding their knowledge in specific domains. By combining ideas from theory-based RL, Bayesian inference, and active exploration, we provide a framework for learning structured, interpretable world models that leverage both the broad knowledge of LLMs and domain-specific experiences without any costly gradient-based learning.
@@ 2.1 Problem statement
We consider an environment framed as a Markov Decision Process $\mathcal{M=(S,A,R,T)}$ with $\mathcal{S}$ the state space, $\mathcal{A}$ the action space, $\mathcal{R:S\times A\times S}\mapsto\mathbb{R}$ the reward function, and $\mathcal{T:S\times A\mapsto S}$ the transition function. In particular, as our main objective is to study world modeling abilities, we consider reward-less environments where no particular task must be solved, i.e., the reward function always returns 0. We focus on textual environments where, given a vocabulary $\mathcal{V}$ and a maximum sentence length $N$, observations $o\in\mathcal{S}\subset\mathcal{V}^{N}$ and actions $a\in\mathcal{A}\subset\mathcal{V}^{N}$ are represented in natural language.
In such a setting, we study how to rapidly learn an LLM-based parametric model $\mathbb{P}^{LLM}$ of the transition function $\mathcal{T}$ (also called forward model) given a budget of experiences $T$ allowed to an agent. This parametric model outputs a probability distribution over all possible next states given a state-action pair. The natural approach from the RL literature to improve this forward model uses gradient-based updates (ha_world_2018; hafner_dream_2020; schrittwieser_mastering_2020). However, the scale of modern LLMs makes such an approach particularly costly. LLMs’ natural in-context learning abilities open up an alternative approach: providing natural language hypotheses in the prompt to improve their predictions. This method is less costly, known as more sample efficient (le_scao_how_2021), and provides human-interpretable updates. Consequently, the parameter we seek to optimize is a set of hypotheses $\mathcal{H}\subseteq\mathcal{V}^{N}$ that, when given as input to the forward model (in addition to state-action pairs), maximizes the likelihood of the experienced next states.
We define the global optimization objective as finding the smallest set of hypotheses $H\subseteq\mathcal{H}$ that maximizes the forward model’s probability of all possible transitions in the environment:
@@ 2.2 WorldLLM
In this section, we present WorldLLM, our methodology for addressing the challenge outlined in the preceding section. Our framework comprises three interconnected components: the Statistician, the Scientist, and the Experimenter. The interactions among these components are illustrated in Figure 1. The Statistician represents our LLM-based forward model (i.e., $\mathbb{P}^{LLM}$) and is used to evaluate the current hypotheses. The set of hypotheses $H$ the Statistician uses is produced by the Scientist, another LLM. These hypotheses are derived from trajectories collected from interactions with the environment using the Experimenter. WorldLLM alternates from during $T$ iterations between evidence collection from the Experimenter and hypotheses refinement from the Scientist. In the subsequent sections, we elaborate on how these modules contribute to our objective’s optimization.
Statistician
We first detail the Statistician module, whose role is to evaluate a set of hypotheses $H_{t}$ on collected transitions. We start by renaming, for the sake of clarity, $\mathbb{P}^{LLM}$ by $\mathbb{P}^{St}$. As described above, the Statistician is a forward model computing the likelihood $\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ of the next state $s^{\prime}$ given a state-action pair $(s,a)$ and a set of hypotheses $H_{t}$. This likelihood is obtained by the LLM by giving it $s$, $a$, and $H_{t}$ in the prompt and taking the probability of $s^{\prime}$ to follow this prompt.
Scientist
To improve the Statistician’s prediction as a forward model, the Scientist’s role is to search through the space of possible sets of hypotheses. To efficiently explore this space, we take inspiration from (wang_hypothesis_2023; piriyakulkij_doing_2024) and use a Bayesian inference framework. In particular, we rely on another LLM $\mathbb{P}^{Sc}$ as the proposal distribution which generates a candidate set of hypotheses following the distribution $\mathbb{P}^{Sc}(\hat{H}_{t}^{i}|\mathcal{D}_{t},\hat{H}_{t})$ given a data set of collected transitions $\mathcal{D}_{t}$ and the current best hypotheses candidate $\hat{H_{t}}$. We assume a uniform prior (i.e., no particular prior) over the space of hypotheses. For this paper, we chose the Metropolis algorithm, a variant of the Metropolis-Hastings algorithm, which iteratively generates candidate hypotheses for $n_{s}teps$. The full algorithm is displayed in Algorithm 1 and more details are provided in Appendix B.
@@
13:     Update score current_score=scorecurrent\_score=score
@@
Experimenter
Given a set of hypotheses produced by the Scientist, the Experimenter must collect transitions that will help assess and refine the hypotheses (i.e., transitions that have a low likelihood under the Statistician). We propose to explore two types of approaches. First, we design oracle agents, allowing us to disentangle the efficiency of data collection and hypotheses refinement in WorldLLM.
For the oracles, we propose the following hard-coded policies:
@@
Then, we introduce RL-based Experimenters. We propose to study three curiosity-based intrinsic rewards derived from $\mathbb{P}^{St}$ (oudeyer_what_2007). First, we introduce RL-LogP, where the Experimenter uses the prediction error (schmidhuber_curious_1991; pathak_curiosity-driven_2017; burda_large-scale_2018) as reward by receiving $r=-log\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ for each collected transition $(s,a,s^{\prime})$. While usually efficient, this method is also known to suffer in stochastic environments where it fails to separate epistemic from aleatoric uncertainty (burda_exploration_2018). Second, we propose RL-ALP, which leverages information gain over the prediction error (a form of Learning Progress) (oudeyer_what_2007; lopes_exploration_2012; schmidhuber_curious_1991) and rewards transitions on which the forward model improves. In particular, we use the absolute value of Learning Progress, which tracks both progress and forgetting in the forward model (oudeyer_what_2007; baranes_r-iac_2009):
@@
Finally, directly tracking Learning Progress at the transition level often leads to noisy estimations. We also propose RL-ALPEXP, a variant of RL-ALP which partitions the transition space into subspaces and computes ALP over each subspace, stabilizing the estimations. While existing works studied how to automatically infer such partitions (baranes_active_2013), we here rely on hand-designed ones, which correspond to the partitions used by oracles. Further details are provided in Appendix A.
@@ 3.1 Experimental setup
Figure 2: Our experimental setup. We use the Playground-Text environment that features objects the agent can interact with, along with textual observations and actions (a). We focus on 4 object types from Playground-Text: Water, Plants, Small Herbivores, and Big Herbivores. We show in (b) the technology tree for combining objects.
To evaluate WorldLLM’s efficiency, we use the Playground-Text environment from (gaven2024sacglamimprovingonlinerl), a textual environment which produces textual observations and receives textual commands. This environment generates a room that contains multiple objects the agent can interact with by grasping or releasing them. The agent can combine two objects, if an interaction is possible, by releasing one onto the other, as illustrated in Figure 2 b. For instance, releasing water on a carrot seed transforms them into grown carrot. In our experiments, we only consider 4 types of objects: Water, Plants (seed or grown) that require water to grow, Small Herbivores (baby or grown) that require a grown plant to grow, and Big Herbivores (baby or grown) that require 2 grown plants to grow. We evaluate WorldLLM’s ability to discover, predict, and explain how to grow Plants, Small Herbivores, and Big Herbivores. We provide more details on the environment in Appendix A. For analysis, we gather the possible transitions into 6 different types: Standing when moving on an object, Holding 1 when grasping an object, Holding 2 when grasping two objects at the same time, Grow Plant when releasing water on a seed, Grow S. Herbivore when releasing a grown plant on a baby small herbivore and Grow B. Herbivore when releasing two grown plants on a baby big herbivore. These types are the ones used by RL-ALPEXP.
@@
We use Phi-3-mini-4k-Instruct [^1] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct for both the Statistician and the Scientist. For our Experimenter, we either use one of our hard-coded oracles or an RL agent. Regarding oracles, O-Ideal collects transitions that match the distribution of the above test set. O-Hardest performs the optimal policy to grow the Small Herbivore and the Big Herbivore accessible. O-Curriculum plays the optimal policy to grow a Plant up to 133 iterations, then plays the optimal policy to grow the accessible Small Herbivore up to 266 iterations, and finally grows the Small Herbivore and the Big Herbivore until the end. See Appendix B.3.1 for more details on the oracles. Concerning the RL-based Experimenters, while it would appear natural to use the LLM already leveraged in the Statistician and Scientist as the policy (e.g., with GLAM), we propose a simpler setup to focus on world modeling abilities. We use a Multi-Layer Perceptron with $2$ hidden layers of $64$ units as the policy, which uses a symbolic representation of Playground-Text’s state as observation. In all experiments below, unless specified otherwise, we use $8$ different random seeds. For each seed, we perform $400$ iterations of the framework (i.e., $T$) where $150$ transitions are collected per iteration and $n_{steps}=5$ steps of the Metropolis algorithm are performed by the Scientist at each iteration. When the Experimenter is an RL agent, we collect $3600$ transitions per iteration to train it and use PPO to update the policy, while only keeping the last $150$ transitions for the Scientist to match the oracles’ hyperparameters. Our test set $\mathcal{D}_{test}$ is composed of 120 Standing, 20 Holding 1, 7 Holding 2, 12 Grow Plant, 6 Grow S. Herbivore and 3 Grow B. Herbivore.
Figure 3: Evolution of log-likelihoods computed by the Statistician (using the Scientist’s last accepted hypotheses) on the test set throughout WorldLLM’s iterations when using different oracles as Experimenters. We average per transition type over 88 random seeds with the standard deviation represented by the error bars. We show at iteration 0 the log-likelihood from the Statistician without any hypotheses given.